OpenStudy (anonymous):

Write a linear equation in slope-intercept form to model the situation. An icicle is 12 inches long and melts at a rate of 1/4 inch per hour. PLEASE HELP

OpenStudy (experimentx):

y = 12-1/4*x (thought not so sure)

OpenStudy (anonymous):

Y = length x = time in hours Y = 12-1/4x

OpenStudy (anonymous):

y=mx+c y=-.25x+12